OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each f(x) = 5(x + 7)^2(x - 7)^3

OpenStudy (anonymous):

Set \[f(x)=5(x+7)^{2}(x-7)^{3}=5(x+7)(x+7)(x-7)(x-7)(x-7)=0\]so a product of factors equaling zero implies one or more factors are zero. \[x+7=0\]\[x=-7\]for TWO of them\[x-7=0\]\[x=7\]for THREE of them. BTW \[5\neq 0\]

OpenStudy (anonymous):

how about the multiplicity though?

OpenStudy (anonymous):

Yep, that's what I meant by TWO and THREE ... ;-)

OpenStudy (anonymous):

yea but where do they go?

OpenStudy (anonymous):

nvm thanks!

OpenStudy (anonymous):

Well, you'd just say something like x = -7 is a zero of f(x) twice repeated, and x= 7 is a zero of f(x) thrice repeated. I guess ...

OpenStudy (anonymous):

Hey can you help me with one more problem?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Find a cubic function with the given zeros. Sqrt 6 , -Sqrt 6, -3

OpenStudy (anonymous):

OK, cubic means to the third power in the independent variable, say x. Begin with \[x = 6^{1/2}\]so\[x-6^{1/2}=0\]then\[x=-6^{1/2}\]so\[x+6^{1/2}=0\]finally\[x=-3\]so\[(x+3)=0\]now combine as a product : \[g(x)=(x+6^{1/2})(x-6^{1/2})(x+3)\]which could be expanded out, but it answers the question as is.

OpenStudy (anonymous):

Note there are others that will fit the given pattern of zeroes, any multiple of g(x) say\[h(x)=5g(x)=5(x+6^{1/2})(x-6^{1/2})(x+3)\]which is what I was hinting at when I said that \[5\neq 0\]because that factor does not alter the function's zeroes.

OpenStudy (anonymous):

but its not one of the choices I have here

OpenStudy (anonymous):

f(x) = x3 - 3x2 - 6x - 18 f(x) = x3 + 3x2 - 6x - 18 f(x) = x3 + 3x2 + 6x - 18 f(x) = x3 + 3x2 - 6x + 18 these are my choices..

OpenStudy (anonymous):

What have you got?

OpenStudy (anonymous):

Ah, we have to multiply the factors out:\[g(x)=(x-6^{1/2})(x+6^{1/2})(x+3)\]\[=(x^2-6^{1/2}x+6^{1/2}x-6)(x+3)\]\[=(x^2-6)(x+3)=x^{3}+3x^{2}-6x -18\]which is the second choice.

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

Find the remainder when f(x) is divided by (x - k) f(x) = 4x3 - 5x2 + 2x + 11; k= 3

OpenStudy (anonymous):

these are the choices 68 136 80 24

OpenStudy (anonymous):

Yo, examine \[p(x)=\frac{4x^3-5x^2+2x+11}{x-3}\]are you familiar with dividing one polynomial into another ?

OpenStudy (anonymous):

no but please I need the answer

OpenStudy (anonymous):

and yes its long division right?

OpenStudy (anonymous):

Yes it is, but rather difficult to compactly describe here though, I'll try ...

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

almost done?

OpenStudy (anonymous):

\[\frac{4x^3-5x^{2}+2x+11}{x-3}\]\[=\frac{4x^3-12x^2+7x^{2}+2x+11}{x-3}\]\[=\frac{4x^{2}(x-3)+7x^2+2x+11}{x-3}=4x^{2}+\frac{7x^2+2x+11}{x-3}\]\[=4x^3+\frac{7x^{2}-21x+23x+11}{x-3}=4x^3+\frac{7x(x-3)+23x+11}{x-3}\]\[=4x^3+7x+\frac{23x+11}{x-3}=4x^3+7x+\frac{23x-69+69+11}{x-3}\]\[=4x^3+7x+\frac{23(x-3)+69+11}{x-3}=4x^3+7x+23+\frac{80}{x-3}\]which we cannot divide further so 80 is the remainder.